OpenStudy (loser66):

Show that if a and b are positive integers and a^3 | b^2, then a| b Please, help

OpenStudy (freckles):

have you thought about using fundamental theorem of arithemetic maybe?

OpenStudy (loser66):

I don't see the result from it. I tried to use \(Let ~ a =\prod p_i^{a_i}\) \(a^3 = \prod p_i^{3a_i}\) Same as b \(b^2=\prod p_i^{2b_i}\) (2,3)= 1, hence if \(a^3 | b^2\) only one case can happen that is \(a_i | b_i\), hence a|b But I don't feel positive on this solution.

OpenStudy (blurbendy):

I think you will need to use induction. You would be able to prove this, but the question says show...

jimthompson5910 (jim_thompson5910):

What you have a^3 | b^2 which means k*a^3 = b^2 where k is an integer What you want to show a | b, in other words m*a = b where m is an integer Hint: k*a^3 = b^2 k*a*a^2 = b^2 (k*a)*(a^2) = b^2 q*(a^2) = b^2 ... let q = k*a

OpenStudy (blurbendy):

@jim_thompson5910 does k need to be positive also?

OpenStudy (loser66):

You meant: \(ka^3 = b^2 \\(ka)a^2 = b*b\) either ka | b or a^2 | b If ka |b , we are done if a^2 | b, then?? \(b = ma^2 = ma* a \) and we done , right?

jimthompson5910 (jim_thompson5910):

a,b are positive b^2, a^3 are positive if k is negative, then k*a^3 is negative which contradicts the fact that b^2 is positive so yes, k has to be positive

OpenStudy (blurbendy):

okay

OpenStudy (loser66):

Still not get @jim_thompson5910 please, guide me more

jimthompson5910 (jim_thompson5910):

The RHS is a perfect square and an integer so the LHS must also be a perfect square and an integer From these steps q*(a^2) = b^2 q = (b^2)/(a^2) q = (b/a)^2 we can see q is a perfect square

OpenStudy (loser66):

Can I do: a^3 | b a^3 | b*b then a^3 | b b = ka^3 = aka^2 a|b is it right?

jimthompson5910 (jim_thompson5910):

if q is a perfect square, then q = m^2 for some integer m q*(a^2) = b^2 ... let q = k*a (m^2)*(a^2) = b^2 ... let m^2 = q

OpenStudy (loser66):

Got you. Thanks a lot.

jimthompson5910 (jim_thompson5910):

no problem